At a family reunion 3 cousins were comparing there ages. Jennifer is 17 years younger than Rene and Rene is 10 years older then Melissa. Their ages total 60 years. How old is Rene?

Answer by Earlsdon(6294)

Let J = Jennifer's age, R = Rene's age, and M = Melissa's age.
J = R-17 "Jennifer is 17 years younger than Rene."
R = M+10 "Rene is 10 years older than Melissa."
J+R+M = 60 "Their ages total 60 years."
Find Rene's age.
Start with:
J+R+M = 60 Now substitute J = R-17 from the first equation.
(R-17)+R+M = 60 Now substitute M = R-10 from the second equation.
R-17+R+R-10 = 60 Combine like-terms.
3R-27 = 60 Add 27 to both sides.
3R = 87 Divide both sides by 3.
R = 29
Rene is 29 years old.